The average horsepower required to raise a 150 kg drum to a height of 20 m in a time of 1 minute is.



a. 0.5 Hp


b. 1 Hp


c. 0.7 Hp


d. 1.2 Hp



1
Expert's answer
2022-02-28T10:43:06-0500
"P=\frac{mhg}{t}=\frac{150*20*9.8}{60}=490\:\rm W=0.7\: hp"

Answer: c